# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ 3.3. The Subfield Metric
(114) $I^{\lambda}_{0}:=\left\{{0}\right\},\quad I^{\lambda}_{1}:=\mathbb{F}_{p}^{\times},\quad I^{\lambda}_{2}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)=0}\right\}\quad\text{and}\quad I^{\lambda}_{3}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)\neq 0}\right\}.$
@@
We call $\mathcal{I}^{\lambda}$ the subfield-trace partition of $\mathbb{F}_{p^{r}}$. Recall that the trace function is uniformly distributed among $\mathbb{F}_{p}$ and that for each $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=mx\mod p$. Hence, the following is a simple consequence.
@@
To extend the partition to $\mathbb{F}_{p^{r}}^{n}$ we first define the subfield-trace decomposition of an element $x\in\mathbb{F}_{p^{r}}^{n}$.
Definition 35.
Given a finite field $\mathbb{F}_{p^{r}}^{n}$ and an element $x\in\mathbb{F}_{p^{r}}^{n}$. The subfield-trace decomposition of $x$ is a $4$-tuple $\pi^{\mathsf{\lambda}}(x)=(\pi^{\mathsf{\lambda}}_{0}(x),\pi^{\mathsf{\lambda}}_{1}(x),\pi^{\mathsf{\lambda}}_{2}(x),\pi^{\mathsf{\lambda}}_{3}(x))$ where, for each $i\in\left\{{0,\ldots,3}\right\}$, we have
(115) $\pi^{\mathsf{\lambda}}_{i}(x)=\left|\,{\left\{{j=1,\ldots,n\;|\;x_{j}\in I^{\lambda}_{i}}\right\}}\,\right|.$
Similar to the Lee and homogeneous metric, we denote the set of all subfield-trace decompositions in $\mathbb{F}_{p^{r}}^{n}$ by
(116) $\mathbb{D}^{\lambda}_{p,r,1,n}=\left\{{\pi\in\left\{{0,\ldots,n}\right\}^{4}\;\big{|}\;\sum_{i=0}^{3}\pi_{i}=n}\right\}.$
Given that $\mathbb{D}^{\lambda}_{p,r,1,n}=\left\{{\pi^{(0)},\ldots,\pi^{(D-1)}}\right\}$, this induces a subfield-trace partition $\mathcal{P}^{\lambda}$ of $\mathbb{F}_{p^{r}}^{n}$ into the vectors of the same subfield-trace decomposition, i.e.,
(117) $\mathcal{P}^{\lambda}:=P^{\lambda}_{\pi^{(0)}}\;|\;\ldots\;|\;P^{\lambda}_{\pi^{(D-1)}}.$
Hence, for a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$, the number of codewords $c\in\mathcal{C}$ of subfield-trace decomposition $\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}$ is given by
(118) $\mathcal{D}_{\pi}^{\lambda}(\mathcal{C}):=\left|\,{P^{\lambda}_{\pi}\cap\mathcal{C}}\,\right|.$
Note also here that every codeword $c\in\mathcal{C}$ with $c\in P^{\lambda}_{\pi}$ has the same subfield weight which is calculated by $\operatorname{wt}_{\lambda}(c)=\pi_{1}+(\pi_{2}+\pi_{3})\lambda$. Thus, the subfield-weight enumerator $\lambda\mathsf{dec}_{\mathcal{C}}$ of a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ can immediately be derived from the subfield-trace enumerator of $\mathcal{C}$.
We are now able to state the MacWilliams identities for the subfield-trace enumerator.
Theorem 36 (MacWilliams-type Identities for the Subfield-Trace Enumerator).
Consider a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given $\rho\in\mathbb{D}^{\lambda}_{p,r,1,n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of subfield-trace decomposition $\rho$ is given by
(119) $\mathcal{D}_{\rho}^{\lambda}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\lambda}_{p,r,1,n}}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C}).$
Here $K^{\lambda}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$, which for a given $p$-th root of unity $\xi$ is defined as
(120) $K^{\lambda}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)}\left(\prod_{i=0}^{3}\binom{\pi_{i}}{t_{i0},\ldots,t_{i3}}\right)\cdot\kappa,$
where
(121) $\kappa=\begin{cases}(p-1)^{t_{01}+t_{11}+t_{21}+t_{03}}(-1)^{t_{31}}(-p)^{t_{22}}(p^{r-1}-p)^{t_{02}+t_{12}}(-p^{r-1})^{t_{13}}(p^{r-1})^{t_{03}}\mathbbm{1}_{t_{23}=t_{32}=t_{33}=0}&\text{if }r\;|\;p,\\ (-1)^{t_{11}+t_{22}+t_{33}+t_{32}}(p-1)^{t_{01}+t_{03}}(1-p)^{t_{2}3}(p^{r-1}-1)^{t_{02}+t_{03}+t_{12}+t_{21}}(1-p^{r-1})^{t_{13}}&\text{otherwise}.\end{cases}$
@@
The multinomial coefficient appearing in the expression for the Krawtchouk coefficient is derived as in the proof of Theorem 30. Thus, it suffices to show that, for any $i\in\left\{{0,\ldots,3}\right\}$, $j\in\left\{{0,\ldots,3}\right\}$ and $a,b\in I^{\lambda}_{j}$ with $a\neq b$ it holds that
(122) $\sum_{x\in I^{\lambda}_{i}}\xi^{\operatorname{Tr}(ax)}=\sum_{x\in I^{\lambda}_{i}}\xi^{\operatorname{Tr}(bx)}.$
That is, in accordance with Example 11, number 3., we want to show
(123) $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)=\sum_{x\in I^{\lambda}_{i}}\chi_{b}(x).$
Therefore, we compute $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)$ in each case and we show that the sum is independent on the choice of $a\in I^{\lambda}_{i}$. For $a=0$ and for every $i\in\left\{{0,\ldots,3}\right\}$ we observe that
(124) $\sum_{x\in I^{\lambda}_{i}}\chi_{a}(x)=|I^{\lambda}_{i}|,$
where the values of $\left|\,{I^{\lambda}_{i}}\,\right|$, for $i=2,3$, can be found in Corollary 34. Similarly, for any $j\in\left\{{0,\ldots,3}\right\}$ and $a\in I^{\lambda}_{j}$ we have
@@
For the remaining cases, and we make use of Lemma 13 and 14 and we identify whether the character $\chi_{a}(x)$ is generating for the subgroup $\mathbb{F}_{p}\subset\mathbb{F}_{p^{r}}$. Recall that, a character $\chi_{a}$ of $\mathbb{F}_{p^{r}}$ is generating for the subgroup $H$ if its kernel is contained in $H$.
Consider first $a\in I^{\lambda}_{1}$, that is $a\in\mathbb{F}_{p}^{\times}$. By $\mathbb{F}_{p}$-linearity of the trace function, we have $\operatorname{Tr}(ax)=a\operatorname{Tr}(x)$ for any $x\in\mathbb{F}_{p^{r}}$. If $r\;|\;p$, for any $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=0$, i.e., $\chi_{a}$ is generating for $\mathbb{F}_{p}$. Hence, by Lemma 13, $\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=p$. Therefore,
@@
Let $a\in I^{\lambda}_{2}$, i.e., we have $\operatorname{Tr}(a)=0$. From the $\mathbb{F}_{p}$-linearity of the trace we obtain that $\operatorname{Tr}(ax)=\operatorname{Tr}(a)x=0$ for any $a\in I^{\lambda}_{2}$ and hence, $\chi_{a}$ is generating for $\mathbb{F}_{p}$ for any $a$ Then, by Lemma 13, we get that
@@
(130) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=0-p=-p&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=0-1=-1&\text{otherwise}.\end{cases}$
Finally, we assume that $a\in I^{\lambda}_{3}$. Note that, if $x=1\in\mathbb{F}_{p}$, then $\operatorname{Tr}(ax)=\operatorname{Tr}(a)\neq 0$. Hence, $\chi_{a}$ is non-generating for $\mathbb{F}_{p}$ over $\mathbb{F}_{p^{r}}$ which implies (using Corollary 14) that
@@
(132) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=0-0=0&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=0-1=-1&\text{otherwise}.\end{cases}$
@@
00 $1$ $2$ $3$
00 $1$ $1$ $1$ $1$
$1$ $p-1$ $p-1$ $p-1$ $-1$
$2$ $p^{r-1}-p$ $p^{r-1}-p$ $-p$ 00
$3$ $p^{r-1}(p-1)$ $-p^{r-1}$ 00 00
(a) Values of $\sum_{x\in I^{\lambda}_{j}}\chi_{a}$ for $a\in I^{\lambda}_{i}$, $i,j\in\left\{{0,\ldots,3}\right\}$ and $r\;|\;p$.
00 $1$ $2$ $3$
00 $1$ $1$ $1$ $1$
$1$ $p-1$ $-1$ $p-1$ $-1$
$2$ $p^{r-1}-1$ $p^{r-1}-1$ $-1$ $-1$
$3$ $(p^{r-1}-1)(p-1)$ $-(p^{r-1}-1)$ $1-p$ $1$
@@ 00 $1$ $1$ $1$ $1$
Let us give an example to illustrate the calculations.

### v3

@@ 3.3. Symmetrized Subfield-Trace Partition
$\ignorespaces\ignorespaces I^{\lambda}_{0}:=\left\{{0}\right\},\quad I^{\lambda}_{1}:=\mathbb{F}_{p}^{\times},\quad I^{\lambda}_{2}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)=0}\right\}\quad\text{and}$
$I^{\lambda}_{3}:=\left\{{x\in\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}\;|\;\operatorname{Tr}(x)\neq 0}\right\}.$
@@
We call $\mathcal{P}^{\lambda}$ the subfield-trace partition of $\mathbb{F}_{p^{r}}$. Recall that for each $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=rx\mod p$ and that the trace function is uniformly distributed among $\mathbb{F}_{p}$. Hence, the following is a simple consequence.
@@
Let $\mathbb{D}^{\lambda}_{n}=\left\{{\pi^{(0)},\ldots,\pi^{(D-1)}}\right\}$ denote the set of $\mathcal{P}^{\lambda}$-decompositions. We denote the induced symmetrized subfield-trace partition $\mathcal{P}_{\mathrm{sym},n}^{\lambda}$ of $\mathbb{F}_{p^{r}}^{n}$ by
(66) $\mathcal{P}_{\mathrm{sym},n}^{\lambda}:=P^{\lambda}_{\pi^{(0)}}\;|\;\cdots\;|\;P^{\lambda}_{\pi^{(D-1)}}.$
Note also here that every codeword $c\in\mathcal{C}$ with $c\in P^{\lambda}_{\pi}$ has the same subfield weight which is calculated by $\operatorname{wt}_{\lambda}(c)=\pi_{1}+(\pi_{2}+\pi_{3})\lambda$. Thus, the subfield-weight enumerator $\lambda\mathsf{dec}_{\mathcal{C}}$ of a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ can immediately be derived from its subfield-trace enumerator.
Proposition 32.
The symmetrized subfield-trace partition $\mathcal{P}_{\mathrm{sym},n}^{\lambda}$ is Fourier-reflexive.
@@
Similar to the proof of Proposition 27 we show that the subfield-trace partition $\mathcal{P}^{\lambda}$ is Fourier-reflexive by calculating the sums $\sum_{x\in I^{\lambda}_{j}}\chi_{a}(x)$, for each $a\in\mathcal{R}$ and each $j\in\left\{{0,\ldots,3}\right\}$.
For $a=0$ and for every $j\in\left\{{0,\ldots,3}\right\}$ we observe that
(67) $\sum_{x\in I^{\lambda}_{j}}\chi_{a}(x)=|I^{\lambda}_{j}|,$
where the values of $\left|\,{I^{\lambda}_{j}}\,\right|$, for $j=2,3$, can be found in Corollary 31. Similarly, for any $i\in\left\{{0,\ldots,3}\right\}$ and $a\in I^{\lambda}_{i}$ we have
@@
For the remaining cases, we make use of Lemmas 12 and 13 and we identify whether the character $\chi_{a}(x)$ is generating for the subgroup $\mathbb{F}_{p}\subset\mathbb{F}_{p^{r}}$. Recall that, a character $\chi_{a}$ of $\mathbb{F}_{p^{r}}$ is generating for the subgroup $H$ if $H$ is not contained in $\ker(\chi_{a})$. Let us first consider $a\in I^{\lambda}_{1}$, that is $a\in\mathbb{F}_{p}^{\times}$. By $\mathbb{F}_{p}$-linearity of the trace function, we have $\operatorname{Tr}(ax)=a\operatorname{Tr}(x)$ for any $x\in\mathbb{F}_{p^{r}}$. If r|pr\;|\;p, for any $x\in\mathbb{F}_{p}$ we have $\operatorname{Tr}(x)=0$, i.e., $\chi_{a}$ is non-generating for $\mathbb{F}_{p}$. Hence, by Lemma 12, $\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=p$. Therefore,
@@
Now, let $a\in I^{\lambda}_{2}$, i.e., we have $\operatorname{Tr}(a)=0$. From the $\mathbb{F}_{p}$-linearity of the trace, we have that $\operatorname{Tr}(ax)=\operatorname{Tr}(a)x=0$ for any $a\in I^{\lambda}_{2},x\in\mathbb{F}_{p}^{\times}$ and hence, $\chi_{a}$ is non-generating for $\mathbb{F}_{p}$ for any $a$ . Then, by Lemma 12, we get that
@@
(73) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=-p&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=-1&\text{otherwise}.\end{cases}$
Finally, we assume that $a\in I^{\lambda}_{3}$. Note that, if $x=1\in\mathbb{F}_{p}$, then $\operatorname{Tr}(ax)=\operatorname{Tr}(a)\neq 0$. Hence, $\chi_{a}$ is generating for $\mathbb{F}_{p}$ over $\mathbb{F}_{p^{r}}$ which implies (using Corollary 13) that
@@
(75) $\sum_{x\in I^{\lambda}_{2}}\chi_{a}(x)=\begin{cases}\sum_{x\in I^{\lambda}_{2}\cup\mathbb{F}_{p}}\chi_{a}(x)-\sum_{x\in\mathbb{F}_{p}}\chi_{a}(x)=0&\text{if }r\;|\;p,\\ \sum_{x\in I^{\lambda}_{2}\cup\left\{{0}\right\}}\chi_{a}(x)-\chi_{a}(0)=-1&\text{otherwise}.\end{cases}$
@@
0 11 22 33
0 11 11 11 11
11 p1p-1 p1p-1 p1p-1 1-1
22 $p^{r-1}-p$ $^{r-1}-p p-p$ 0
33 $p^{r-1}(p-1)$ $p^{r-1}$ 0 0
Table 3. Values of $\sum_{x\in I^{\lambda}_{j}}\chi_{a}$ for $a\in I^{\lambda}_{i}$, $i,j\in\left\{{0,\ldots,3}\right\}$ and r|pr\;|\;p.
0 11 22 33
0 11 11 11 11
11 p1p-1 1-1 p1p-1 1-1
22 $p^{r-1}-1$ $p^{r-1}-1$ 1-1 1-1
33 $(p^{r-1}-1)(p-1)$ $(p^{r-1}-1)$ 1p1-p 11
@@
For some cases, this partition is induced by partitions of an orbit. However, as we could not verify this for all instances, we leave it as an open question.
We are now able to state the MacWilliams identities for the subfield-trace enumerator whose existence is given by [18, Theorem 3.5].
Corollary 33 (MacWilliams-type Identities for the Symmetrized Subfield-Trace Enumerator).
Consider a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given $\rho\in\mathbb{D}^{\lambda}_{n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of subfield-trace decomposition $\rho$ is given by
(76) $\mathcal{D}_{\rho}^{\lambda}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\lambda}_{n}}K^{\lambda}_{\pi}(\rho)\mathcal{D}_{\pi}^{\lambda}(\mathcal{C}).$
Here $K^{\lambda}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$ which, for a given $p$-th root of unity $\xi$, is defined as
(77) $K^{\lambda}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)}\left(\prod_{i=0}^{3}\binom{\pi_{i}}{t_{i0},\ldots,t_{i3}}\right)\cdot\kappa^{\lambda},$
where $\kappa^{\lambda}$ takes the values
[…]
$(p-1)^{t_{01}+t_{03}}(-1)^{t_{11}+t_{22}+t_{33}+t_{32}}(1-p)^{t_{23}}(p^{r-1}-1)^{t_{02}+t_{03}+t_{12}+t_{21}}(1-p^{r-1})^{t_{13}},$
if r|pr\;|\;p and $r\not\;\mid p$, respectively.
Proof.
The multinomial appearing in the expression of the Krawtchouk coefficient $K^{\lambda}_{\pi}(\rho)$ is derived as in the proof of Corollary 28. The formula follows using the values of Tables 3 or 4, respectively, with the multiplicities provided by the composition $t\in\mathrm{Comp}^{\lambda}_{\pi}(\rho)$. ∎
An illustrative example can be found in A.3.
Lastly, we want to highlight that the symmetrized subfield-trace partition (see (66)) is indeed the coarsest partition that allows for the MacWilliams identities to exist and to recover the subfield weight enumerator. The proof follows similar arguments to the proof of Lemma 29 and is therefore omitted.
Lemma 34.
Given a finite field $\mathbb{F}_{p^{r}}$ and a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$. The symmetrized subfield-trace partition is the coarsest symmetrized partition which is Fourier-reflexive and allows to recover the subfield-weight enumerator of $\mathcal{C}$.
4. Linear Programming Bound
Given a finite chain ring $\mathcal{R}$, an additive weight $\mathrm{wt}:\mathcal{R}^{n}\to\mathbb{Q}$ and an $\mathcal{R}$-linear code $\mathcal{C}$, recall the definition of the decomposition enumerator of $\mathcal{C}$
with respect to the weight $\operatorname{wt}$. Moreover, if $\left|\,{\mathbb{D}^{\operatorname{wt}}_{n}}\,\right|=D$, we denote by $\pi^{(0)}=(n,0,\dots,0)$ the decomposition of the all-zero vector, and all the other decompositions in $\mathbb{D}^{\operatorname{wt}}_{n}$ by $\pi^{(1)},\dots,\pi^{(D-1)}$.
The MacWilliams-type identities for the Lee decomposition enumerator, given in Corollary 24, allow us to straightforwardly derive an LP bound for any $\mathcal{R}$-linear code and any additive weight. Moreover, the MacWilliams-type identities given in Corollary˜28 and Corollary˜33 allow to derive an LP bound for the homogeneous and subfield metric respectively, which is less complex (i.e., that involves fewer variables).
From now on, $B_{\mathcal{R}}(n,\operatorname{d}_{\operatorname{wt}})$ denotes the maximum number of codewords of an $\mathcal{R}$-linear code of length $n$ and minimum $\mathrm{wt}$-distance $\operatorname{d}_{\operatorname{wt}}$.
Theorem 35 (Linear Programming Bound).
For a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ of minimum distance $\operatorname{d}_{\operatorname{wt}}$ it holds that
(79) $B_{\mathcal{R}}(n,\operatorname{d}_{\operatorname{wt}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\right),$
  • $\mathcal{D}_{\pi^{(0)}}^{\operatorname{wt}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})\geq 0$ for all $1\leq i\leq D-1$ and $\mathcal{D}_{\pi^{(i)}}^{\operatorname{wt}}(\mathcal{C})=0$ for all $i\in I$;
where $I=\{i\ \;|\;\ x\in\mathcal{R}^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\mathrm{wt}(x)<\operatorname{d}_{\operatorname{wt}}\}$.
If $\operatorname{wt}$ is an additive weight of $\mathcal{R}^{n}$ and $\mathcal{D}_{\pi}^{\operatorname{wt}}(\mathcal{C})$ is the Lee partition enumerator, the LP bound requires a large number of variables. Thus, the program cannot run efficiently and is not computationally feasible.
4.1. Lee metric
The Lee scheme, introduced in [23] and investigated in [38, 39, 24], provides a framework for studying the maximum size of a linear Lee-metric code. In the following we will refine the Lee-metric LP bound Theorem˜35.
Throughout this section, let $\mathcal{R}=(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ and let $M$ be the maximum value of the Lee-weight in $\mathbb{Z}/{p^{s}}\mathbb{Z}$. Given a $\mathbb{Z}/{p^{s}}\mathbb{Z}$-linear code $\mathcal{C}$ of length $n$, let $c\in\mathcal{C}$ denote a codeword. In the Hamming metric, whenever we multiply each component of $c$ by a unit $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$, we obtain a codeword $r\cdot c$ of the same Hamming weight as $c$. This property does not hold in the Lee metric. However, the Lee decompositions of $c$ and $r\cdot c$ are related. Indeed, the Lee decomposition of $r\cdot c$ is a permutation of the Lee decomposition of $c$. That is, if $\pi^{\mathsf{L}}(c)=(\pi^{\mathsf{L}}_{0},\dots,\pi^{\mathsf{L}}_{M})$ is the Lee decomposition of $c$, then the Lee decomposition of $r\cdot c$ is given by
where $\sigma_{r}(i)=k$ such that $rk\equiv i\pmod{p^{s}}$.
Lemma 36 ([40, Lemma 1]).
Given a Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{n}$, a unit element $r\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{\times}$ and $\sigma_{r}$ a permutation on the set $\left\{{1,\ldots,M}\right\}$, we define $\sigma_{r}(\pi):=(\pi_{0},\pi_{{\sigma_{r}}(1)},\dots,\pi_{\sigma_{r}(M-1)})$. Then, for any linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$, it holds that $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\sigma(\pi)}^{\mathsf{L}}(\mathcal{C})$.
For any Lee decomposition $\pi\in\mathbb{D}^{\mathsf{L}}_{n}$, we define the set
Since $\mathcal{D}_{\pi}^{\mathsf{L}}(\mathcal{C})=\mathcal{D}_{\rho}^{\mathsf{L}}(\mathcal{C})$ for every $\rho\in S(\pi)$, an additional constraint arises in the formulation of the LP problem for linear Lee-metric codes.
Let $B_{p,1,s,n}(n,\operatorname{d}_{\mathsf{L}})$ denote the maximum number of codewords in a linear code of length $n$ and minimum Lee-distance $\operatorname{d}_{\mathsf{L}}$ over $\mathbb{Z}/{p^{s}}\mathbb{Z}$.
Theorem 37 (Linear Programming Bound for Lee-metric Codes).
For a linear code $\mathcal{C}\subseteq(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}$ of minimum Lee distance $\operatorname{d}_{\mathsf{L}}$ it holds that
(82) $B_{\mathbb{Z}/{p^{s}}\mathbb{Z}}(n,\operatorname{d}_{\mathsf{L}})\leq\max\left(\sum_{i=0}^{D-1}\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\right),$
  • $\mathcal{D}_{\pi^{(0)}}^{\mathsf{L}}(\mathcal{C})=1$;
  • $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})\geq 0$ for all $1\leq i\leq D-1$ and $\mathcal{D}_{\pi^{(i)}}^{\mathsf{L}}(\mathcal{C})=0$ for all $i\in I$;
  • and
where $I=\{i\ \;|\;\ x\in(\mathbb{Z}/{p^{s}}\mathbb{Z})^{n}\text{ such that }\pi(x)=\pi^{(i)}\text{ and }\operatorname{wt}_{\mathsf{L}}(x)<\operatorname{d}_{\mathsf{L}}\}$.
The last constraint is derived from the fact that
5. Conclusion
In this paper, we investigated MacWilliams-type identities for codes over finite chain rings equipped with additive metrics. Our approach is based on studying partitions of the ambient space that refine the corresponding weight partitions while remaining Fourier-reflexive. For the Lee metric, we showed that the partition induced by the action of the full group of linear isometries is already the coarsest Fourier-reflexive refinement of the weight partition. In contrast, for the homogeneous and subfield metric, we identified new symmetrized partitions that are significantly coarser than the orbit partitions induced by subgroups of the linear isometry group, yet remain Fourier-reflexive. These partitions are, moreover, the coarsest symmetrized partitions for which MacWilliams-type identities hold. The presented identities can be used to formulate LP bounds in each of the metrics.
Appendix A Examples
In this section we provide computational examples of the partition-based Mac-Williams identities in the Lee, homogeneous and subfield metric presented in Corollaries 24, 28 and 33, respectively.
A.1. Lee Weight Partition
We start by giving an example of the MacWilliams-type identities for the Lee weight partition given in Corollary 24.